OpenStudy (anonymous):

x^2+4x-32=0 what are the solutions ?

OpenStudy (saifoo.khan):

Factorize!

OpenStudy (saifoo.khan):

need help?

OpenStudy (anonymous):

- Yes , sorry . :/

OpenStudy (saifoo.khan):

- lol, np!

OpenStudy (saifoo.khan):

x^2 +8x - 4x - 32 = 0 Find commons for pairs! x(x + 8) - 4(x + 8) = 0 (x + 8) (x - 4) = 0 now solve further to find solutions: x + 8 = 0 & x -4 = 0 x = -8 & x = 4